# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v6

@@ 4.3 Discussions
Pre-trained knowledge retention and potential pre-trained model improvement. To assess transfer beyond the zero-shot limits of prior works, we evaluated on unseen datasets—CIFAR100 (Krizhevsky et al., 2009), Places365 (Zhou et al., 2017), and ImageNet-1k (Deng et al., 2009) (Table 2). Unlike methods that rely solely on zero-shot predictions (Yu et al., 2024; Xu et al., 2024) and are constrained by CLIP’s baseline, our approach integrates new knowledge into the pre-trained space while preserving existing capabilities. This yields superior zero-shot transfer accuracy, demonstrating that CoDyRA effectively expands general knowledge and improves overall representations.
Rank allocation across different datasets. We visualize the ranks merged into the pre-trained weights for different datasets in Fig. 6 and Fig. 6 (more in Appendix Fig. 11 to Fig. 17). The results show that CoDyRA adapts the number of active ranks differently across locations and tasks, reflecting the varying demands of each dataset. These strike a balance between plasticity and stability, demonstrating the method’s ability to regulate forgetting while enabling effective task learning.
For the Aircraft dataset, we observe that a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig. 6(b)). Additionally, a higher concentration of ranks is assigned to deeper layers, particularly layers 6 to 10 (Fig. 6(c)).
In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig. 6(b)). Additionally, rank allocation is primarily concentrated in the first and last layers (Fig. 6(c)), indicating a distinct adaptation pattern that balances knowledge retention and adaptation across datasets. These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge. We provide more visualizations and analyses in Appendix B.4.
@@ Methods
LWF (Li & Hoiem, 2017) 129.6M None
ZSCL (Zheng et al., 2023) 129.6M None
MoE-Adapters (Yu et al., 2024) 59.8M 13.35M
RAIL (Xu et al., 2024) N/A 24.18M / 9.01M
@@ 5 Conclusion
We introduced CoDyRA, a continual learning framework for CLIP that dynamically selects and minimizes LoRA ranks based on their learned importance. Guided by sparsity-promoting regularization, CoDyRA balances plasticity with stability, without storing past data, task information, or relying on assumptions. By keeping updates close to the model’s previous state while preserving essential ranks, our method mitigates forgetting and enhances representation quality. Extensive experiments across diverse continual learning benchmarks demonstrate that CoDyRA achieves state-of-the-art performance, effectively balancing downstream adaptation and knowledge retention.
Future works. CoDyRA updates the model continually by balancing learning and forgetting in the current task. While enabling a flexible and efficient CL process with strong results, future work will further refine the modeling of rank relationships, inter-task dependencies, and forgetting measures.
Reproducibility Statement
Detailed experiment settings and implementation details are provided in Sec. 4.1 and Appendix B.
@@ Appendix A Additional Details of the Proposed Method
We apply proximal gradient descent (Beck & Teboulle, 2009) to handle the non-differentiable $\ell_{1}$ regularization of the importance weights (Ding et al., 2023). At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
@@
$\mathcal{T}_{\kappa}(x)=\mathbbm{1}(\lvert x\rvert>\kappa)\cdot(x+\mathrm{sign}(x)\cdot\kappa),$ (7)
@@
Appendix B More Experimental Results and Ablations
@@
Further details of the experiment settings. The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010). In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024). In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL. For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
B.1 More Experiment Results
Table 5: Extended results for Table 1, comparing continual fine-tuning with different LoRA settings and the LoRA variant of LAE (Gao et al., 2023).
Method
@@ B.1 More Experiment Results
   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
@@
B.1.1 Comparison to fixed‐rank LoRA baselines.
Table 5 reports results for fixed‐rank LoRA (r = 4, 8, 16) and the LoRA variant of LAE. Three key observations emerge: (1) As LoRA rank increases, Last accuracy improves while Average accuracy declines, aligning with Fig. 4: a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. (2) Fixed‐rank LoRA shows a substantial Last‐accuracy gap compared to our adaptive method. This suggests that without dynamic rank selection, low‐relevance ranks (potentially encoding noise) are retained in the model, impairing generalization to unseen domains and accumulating inter‐task interference over time, resulting in severe performance drop. (3) Leveraging fast online LoRA updates, slow EMA aggregation in offline LoRA, and energy-based logit ensembling, LAE substantially outperforms fixed-rank LoRA (r = 16). Nonetheless, unlike our method, fixed-rank LoRA lacks fine-grained control over the contribution of individual ranks within each module.
@@
Method
Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
@@
   Zero-shot (Radford et al., 2021) 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 56.9
   LwF (Li & Hoiem, 2017) 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
   iCaRL (Rebuffi et al., 2017a) 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
   LwF-VR (Ding et al., 2022) 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
   WiSE-FT (Wortsman et al., 2022) 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
   ZSCL (Zheng et al., 2023) 55.7 50.9 30.4 41.4 49.3 71.8 36.3 77.5 55.3 53.4 53.2
@@
   LwF (Li & Hoiem, 2017) 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
   iCaRL (Rebuffi et al., 2017a) 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
   LwF-VR (Ding et al., 2022) 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
   WiSE-FT (Wortsman et al., 2022) 26.7 86.5 64.4 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
   ZSCL (Zheng et al., 2023) 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
@@
   LwF (Li & Hoiem, 2017) 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
   iCaRL (Rebuffi et al., 2017a) 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
   LwF-VR (Ding et al., 2022) 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
   WiSE-FT (Wortsman et al., 2022) 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
   ZSCL (Zheng et al., 2023) 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
@@
B.1.2 Additional results on full-shot MTIL.
To further extend our evaluation, we report full-shot results on the MTIL benchmark (Table 6). For fairness, we adopt the same domain-prediction design used in recent methods. Consistent with Tables 1 and 3, CoDyRA achieves state-of-the-art performance, surpassing all baselines on both Transfer (unseen-domain generalization) and Last (continual downstream learning).
Table 6 also compares against prompt-based approaches (L2P (Wang et al., 2022e), DualPrompt (Wang et al., 2022d), S-Prompt (Wang et al., 2022c)). Although these methods mitigate task forgetting via key–value matching for prompt selection, they neglect the preservation of pre-trained capabilities—evidenced by substantial drops in transfer accuracy—and further underperform recent domain-prediction approaches on last accuracy.
@@
B.2 Effects of Additional Adaptive Prediction Head on Visual Representations
The related continual learning methods focus on updating the representation with the backbone model (e.g., (Wang et al., 2022e; Zheng et al., 2023; Yu et al., 2024), and CoDyRA) or newly added representation alignment or prediction head (McDonnell et al., 2024; Xu et al., 2024). These two types of methods can be seen as orthogonal. We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in (Xu et al., 2024).
We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 8. Note that the Dual version of RAIL (Xu et al., 2024) stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 3 for a fair comparison.

### v7

@@ 4.2 Discussions
Pre-trained knowledge retention and potential improvement. We assess transfer capability on unseen datasets: CIFAR100 [35], Places365 [100], and ImageNet-1k [12] (Table 3). Unlike prior methods [93, 90] that freeze the backbone and remain bounded by CLIP’s initial zero-shot performance, CoDyRA continually integrates new knowledge into the parameter space. This allows our model to surpass the original zero-shot baseline, demonstrating that CoDyRA effectively expands general knowledge and potentially improves overall representations.
Rank allocation analysis. We visualize rank allocation across datasets in Fig. 6 and Fig. 6 (extended in Appendix Fig. 1218). The results show that CoDyRA adapts the number of active ranks differently across locations and tasks, reflecting the varying demands of each dataset. These strike a balance between plasticity and stability, demonstrating the method’s ability to regulate forgetting while enabling effective task learning.
For Aircraft, ranks concentrate in the text encoder’s Attention Output projection (Fig. 6(b)) and deeper layers 6–10 (Fig. 6(c)). In contrast, for Pets, allocation shifts to the vision encoder (Attn Value/Output and MLP Projection; Fig. 6(b)), primarily targeting the first and last layers (Fig. 6(c)).
These distinct patterns highlight CoDyRA’s ability to dynamically adjust capacity for effective domain adaptation while preserving existing knowledge. Further analysis is provided in Appendix B.8.
@@ Methods
LWF [38] 129.6M None
ZSCL [99] 129.6M None
MoE-Adapters [93] 59.8M 13.35M
RAIL [90] N/A 24.18M / 9.01M
@@ 5 Conclusion
We introduced CoDyRA, a dynamic rank-selective framework for continual learning in large pre-trained models, including both CLIP and LLMs. Guided by sparsity-promoting regularization, CoDyRA automatically identifies the ranks needed, keeping updates close to the model’s previous state while preserving essential ranks. This approach eliminates the need for data replay, task identifiers, or complex architectural expansion. Extensive evaluation across diverse benchmarks confirms that CoDyRA achieves superior performance, effectively balancing learning and knowledge retention across modalities and scales.
Future works. CoDyRA updates the model continually by balancing learning and forgetting in the current task. While enabling a flexible and efficient CL process with strong results, future work will further refine modeling of rank relationships, inter-task dependencies, and forgetting measures.
Impact Statement
CoDyRA’s sparse, adaptive, importance‐aware low‐rank subspace learning significantly reduces compute and memory demands for continual learning, enabling on‐device personalization in domains such as mobile health and education and empowering resource‐constrained users and organizations. As a fundamental research effort, CodyRA does not address potential biases or hazards in training data, and its incremental updates could introduce privacy risks or stealthy backdoors. We therefore recommend transparent update logging, regular audits, and differential‐privacy safeguards. Although more efficient than full fine‐tuning, large-scale deployment still incurs energy costs; developers should monitor resource use, leverage low-power hardware, and explore carbon offsets.
Acknowledgments and Disclosure of Funding
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
@@ Appendix A Additional Details of the Proposed Method
We apply proximal gradient descent [4] to handle the non-differentiable $\ell_{1}$ regularization of the importance weights [14]. At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
@@
$\mathcal{T}_{\kappa}(x)=\mathbbm{1}(\lvert x\rvert>\kappa)\cdot(x{-}\mathrm{sign}(x)\cdot\kappa),$ (7)
@@
A.2 More Discussions on Analytical Results of Sec. 3.2
A.2.1 Takeaway 1 - On Applying LoRA to All Modules
The original LoRA paper applies LoRA only to selected modules (mainly the attention components). Many subsequent LoRA and LoRA-based continual learning works (e.g., InfoLoRA, MoE-Adapters, O-LoRA, etc.) also apply LoRA to only part of the Transformer modules, though the specific choices vary. To clarify this design space, we conducted an analysis to determine which module configuration is most effective. Our results show that applying LoRA to all layers yields clear benefits.
While the theoretical analysis of standard Transformers and full LoRA training dynamics is complex, the empirical Neural Tangent Kernel (eNTK) [50] offers a practical framework for theoretical investigation. The necessity of applying LoRA to all layers, rather than just attention modules, can be explained through the lens of the eNTK .
The eNTK of LoRA approximates that of Full Fine-Tuning (FullFT) only when LoRA is applied to the layers that contain the majority of the parameters. Since the kernel is based on the dot products of gradients ($K(i,j)=g_{i}\cdot g_{j}$), layers with more parameters (e.g., MLPs) typically exert the most influence on the kernel. Restricting updates to attention layers limits the model’s ability to approximate the learning dynamics of FullFT, resulting in "slower learning" or reduced plasticity.
A.2.2 Takeaway 2 - On the Rank-Dependent Plasticity-Stability Balance
We formalize the causality of this balance by defining the trade-off as an informational dilemma rooted in the subspace dimensionality:
High Rank $\rightarrow$ High Plasticity / Low Stability. The rank ($r$) dictates the intrinsic dimensionality (degrees of freedom) of the update subspace $\Delta W$. A high rank maximizes the informational capacity of the parameter shift. In the NTK context, maximizing rank pushes the update closer to the FullFT regime. While this allows for rapid acquisition of new task knowledge (High Plasticity), the larger informational perturbation maximizes disruption to the pre-trained knowledge manifold, leading to catastrophic forgetting (Low Stability).
Low Rank $\rightarrow$ High Stability / Low Plasticity. Conversely, restricting rank constrains the degrees of freedom. This forces the model to encode only the minimal, essential information for the new task. This minimization of total informational perturbation ensures the model remains informationally proximal to its previous state—a necessary condition for stability.
A lower rank $r$ of the LoRA implies a smaller dimension for the update $\Delta W$. A low rank acts as a structural regularizer. It forces the update to be the "simplest" change necessary to solve the current task. This minimizes the norm $\Delta W||_{F}$ and reduces the probability that the update vector on the critical subspace of previous tasks.
A.2.3 Takeaway 3 - On the Necessity of Adaptive Optimization
The need for adaptive rank selection stems from the functional variance of model components and the different information requirements of tasks.
Following the theory on intrinsic dimension of objective landscapes [36], different tasks require different degrees of freedom to be solved. And, not all layers contribute equally to the objectives. Adaptivity allows the model to allocate higher ranks to layers with high gradient influence (e.g., specific MLPs) while enforcing sparsity on less critical layers.
A.3 Connections Between Rank Minimizing and Forgetting
As shown in Eq. (2) and (3), we minimize the rank number to regularize the updating strength $\|\Delta\mathbf{W}^{t,m}\|$, $R(\Delta\mathbf{W}^{t,m})$. Minimizing $\|\mathbf{w}^{t,m}\|_{1}$ with a small rank number encourages minimizing of $\|\Delta\mathbf{W}^{t,m}\|_{F}$ if given $\mathbf{A}^{t,m}$ and $\mathbf{B}^{t,m}$ (according to Eq. (2)). It limits the update $\Delta\mathbf{W}^{t,m}$ to the minimal dimension necessary. It softly reduces capacity in updates conflicting with critical information encoded in the learned parameters, without introducing additional assumptions.
Another line of work regularizes updates (e.g., LoRA weights or gradients) to lie in subspaces orthogonal to existing parameters [80, 40]. Orthogonal-regularization methods and CoDyRA both aim to mitigate interference, but adopt different strategies with distinct trade-offs, which can be summarized along three aspects.
(1) Orthogonal regularization models and aims to constrain updates to lie in subspaces orthogonal to previous parameters, which effectively reduces interference but also limits the ability to acquire new knowledge. This rigid geometric constraint may suppresses learning new knowledge, and thus is typically softened or relaxed in practice. In contrast, CoDyRA regularizes update simplicity (via rank sparsity) rather than geometry, allowing alignment with prior knowledge when beneficial while keeping updates compact to avoid unintended overwriting. Both of them aims to alleviate interference (for stability) but have trade-off for learning new tasks (for plasticity).
(2) Explicit orthogonal-regularization methods typically require storing the information of already the learned subspace (e.g., projection matrices or singular vectors) to compute orthogonality losses, introducing additional memory overhead and dependence on past task statistics. In contrast, CoDyRA mitigates interference implicitly through sparsity over rank importance. This strictly local objective requires no access to previous tasks, aligning with our goal of natural, memory-free PTM updates.
(3) Standard orthogonal methods typically assume a fixed rank and focus on rotating it. When this rank exceeds a task’s intrinsic dimensionality, the extra dimensions behave as noise, causing unnecessary weight drift. CoDyRA actively prunes such dimensions: a rank pruned to zero induces, by definition, “zero interference” with previous tasks. This yields “zero overlap” for irrelevant dimensions while preserving flexible updates for task-relevant ones.
By mitigating forgetting through different mechanisms and trade-offs, CoDyRA and orthogonal-based methods are complementary. For example, the task-relevant ranks selected by CoDyRA (with nonzero ${\textnormal{w}}^{t,m}$) can be further regularized using orthogonal constraints.
Appendix B More Experiments
B.1 Further Details of Experiments.
@@ Appendix B More Experiments
Details of CLIP CL experiment settings. The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft [49], Caltech101 [19], CIFAR100 [35], DTD [10], EuroSAT [25], Flowers [57], Food [6], MNIST [13], OxfordPet [58], Cars [34], and SUN397 [89]. In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 [19] excluded to prevent domain overlap, following the protocol in [90]. In line with [90], we evaluate MTIL under both 5-shot and full-shot settings, while X-TAIL is evaluated using a 16-shot setting.
Details of LLM CL experiment settings. We adopt the TRACE benchmark [81], designed for evaluating continual learning in LLMs such as LLaMA [72, 16] and Gemma [70]. TRACE comprises eight diverse datasets standardized into a unified format, spanning domain-specific understanding, multilingual processing, code generation, and mathematical reasoning. Specifically, the tasks are: C-STANCE [98], FOMC [67], MeetingBank [28], Py150 [47], ScienceQA [46], NumGLUE-cm [55], NumGLUE-ds [55], and 20Minuten [22]. In our setup, each task contributes 5,000 training instances and 2,000 testing instances, totaling 40,000 training and 16,000 testing samples.
Given this diversity, we employ task-specific metrics to capture performance nuances: Accuracy is used for classification and reasoning tasks (C-STANCE, FOMC, ScienceQA, NumGLUE); ROUGE-L for summarization (MeetingBank); SARI for multilingual text simplification (20Minuten); and a similarity score for code generation (Py150).
B.2 Evaluation Metrics
To strictly evaluate the plasticity and stability of our method, we employ the following metrics. Let $N$ denote the total number of learned tasks, and $A_{i,j}$ denote the performance (e.g., accuracy, ROUGE, or similarity score) on task $j$ after training on task $i$.
CLIP Continual Learning Metrics.
For the CLIP experiments (MTIL and X-TAIL), we report the following metrics as in [99, 93, 90]:
  • Transfer Accuracy (Transfer): indicates the model’s zero-shot performance on future domains before they are learned, measures the extent to which the pre-trained zero-shot ability is preserved (or improved) throughout incremental learning.
  • Average Accuracy (Average): indicates the average accuracy of all learning steps across all domains. It captures the comprehensive performance stability throughout the entire incremental training process.
  • Last Accuracy (Last): represent the model’s performance on all seen domains after the training is fully completed.
LLM Continual Learning Metrics.
For the TRACE benchmark, we adopt the standard metrics as in [81]:
  • Overall Performance (OP): The average performance across all learned tasks at the end of training.
    $\text{OP}=\frac{1}{N}\sum_{j=1}^{N}A_{N,j},$ (8)
    where $A_{i,j}$ represents the performance on task $j$ after learning task $i$.
  • Backward Transfer (BWT): Measures the average performance degradation (forgetting) of previous tasks $j<N$ after learning new tasks. A lower BWT indicates less forgetting (better stability).
    $\text{BWT}=\frac{1}{N-1}\sum_{j=1}^{N-1}(A_{j,j}-A_{N,j}).$ (9)
B.3 Implementation Details.
CLIP experiments. Following established protocols [99, 93, 90], we employ the CLIP ViT-B/16 backbone [62]. CoDyRA is applied to all pre-trained weight matrices in both vision and text encoders, initialized with a rank of 16. CoDyRA is optimized for 500 iterations per task using AdamW [45], with dense training ratios of 0.5 for X-TAIL and 0.7 for MTIL, setting $\kappa_{\text{max}}=0.005$.
LLM experiments. We adhere to the variable epoch schedule of prior work [60] ({2, 1, 3, 2, 1, 2, 2, 3} epochs). We use an initial rank of 16, a dense training ratio of 0.5, and set $\kappa_{\text{max}}=0.005$. Optimization employs a learning rate of $1\times 10^{-4}$ and a batch size of 4. Training utilizes DeepSpeed ZeRO-2 with BF16 mixed-precision and a maximum context length of 1024 tokens. All experiments were conducted on a cluster of four Nvidia H100 GPUs.
B.4 More Experiment Results
B.4.1 Cross-domain task-agnostic incremental learning.
Table 5 presents results on the X-TAIL benchmark [90], a more challenging setting where inference is performed over a mixed label space across all domains. CoDyRA outperforms prior methods across all metrics without requiring auxiliary components, surpassing the zero-shot upper bound in transfer accuracy.
Table 5: Comparisons on X-TAIL for each domain. The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [93, 90].
    Method Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
@@ B.4 More Experiment Results
   Zero-shot [62] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [38] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [85] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [64] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [99] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
[…]
   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
[…]
Average
   LwF [38] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [85] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [64] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [99] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
[…]
   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
[…]
Last
   LwF [38] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [85] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [64] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [99] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
[…]
   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
[…]
Table 6: Extended results for Table 5 on X-TAIL for each domain, comparing continual fine-tuning with different LoRA settings and other LoRA variants.
    Method Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
CLIP
   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
Transfer
   Zero-shot [62] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
@@
B.4.2 Comparison to fixed‐rank LoRA baselines.
Table 6 reports results for fixed‐rank LoRA (r = 4, 8, 16) and the LoRA variants, including LAE [20], TreeLoRA [60], InfLoRA [40] and CL-LoRA [24]. Several key observations emerge: (1) As LoRA rank increases, Last accuracy improves while Average accuracy declines, aligning with Fig. 3: a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. (2) Fixed‐rank LoRA shows a substantial Last‐accuracy gap compared to our adaptive method. This suggests that without dynamic rank selection, low‐relevance ranks (potentially encoding noise) are retained in the model, impairing generalization to unseen domains and accumulating inter‐task interference over time, resulting in severe performance drop. (3) Leveraging fast online LoRA updates, slow EMA aggregation in offline LoRA, and energy-based logit ensembling, LAE substantially outperforms fixed-rank LoRA (r = 16). Nonetheless, unlike our method, fixed-rank LoRA lacks fine-grained control over the contribution of individual ranks within each module.
@@
    Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
@@
   Zero-shot [62] 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 56.9
   LwF [38] 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
   iCaRL [63] 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
   LwF-VR [15] 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
   WiSE-FT [85] 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
   ZSCL [99] 55.7 50.9 30.4 41.4 49.3 71.8 36.3 77.5 55.3 53.4 53.2
@@
   LwF [38] 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
   iCaRL [63] 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
   LwF-VR [15] 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
   WiSE-FT [85] 26.7 86.5 64.4 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
   ZSCL [99] 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
@@
   LwF [38] 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
   iCaRL [63] 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
   LwF-VR [15] 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
   WiSE-FT [85] 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
   ZSCL [99] 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
@@
B.4.3 Additional results on full-shot MTIL.
To further extend our evaluation, we report full-shot results on the MTIL benchmark (Table 7). For fairness, we adopt the same domain-prediction design used in recent methods. Consistent with Tables 5 and 1, CoDyRA achieves state-of-the-art performance, surpassing all baselines on both Transfer (unseen-domain generalization) and Last (continual downstream learning).
Table 7 also compares against prompt-based approaches (L2P [84], DualPrompt [83], S-Prompt [82]). Although these methods mitigate task forgetting via key–value matching for prompt selection, they neglect the preservation of pre-trained capabilities—evidenced by substantial drops in transfer accuracy—and further underperform recent domain-prediction approaches on last accuracy.
@@
B.4.5 Analysis in forgetting
We provide a dedicated forgetting analysis in Table 9, quantifying forgetting via Backward Transfer (BWT). It reports BWT of CoDyRA and some relevant CL methods with representation updating. Fixed-rank LoRA serves as a baseline and exhibits significant forgetting due to unconstrained updates without regularization. ZSCL relies on the replay of reference data to help mitigate forgetting. MoE-Adapters isolates task-wise knowledge within specific adapters and adopts domain prediction to further reduce interference. InfLoRA constrains parameter updates to a direction orthogonal to prior tasks to reduce interference. CoDyRA effectively mitigates forgetting by minimizing the rank of updates via sparsity regularization. This ensures the model learns effectively while remaining close to its original state, demonstrating strong retention capabilities with a BWT of only $1.87\%$.
Moreover, in Table 2, we validate our method in the language domain. The low BWT scores observed in this setting further support our claims of reduced forgetting across architectures.
Table 9: Forgetting analysis. We report Backward Transfer (BWT) averaged across all X-TAIL tasks, where lower values indicate better retention.
Methods ZSCL MoE-Adapter InfLoRA LoRA (rr=4) LoRA (rr=8) LoRA (rr=16) CoDyRA
BWT (%) $\downarrow$ 8.78 5.43 3.12 13.94 14.24 14.86 1.87
B.5 Effects of Additional Adaptive Prediction Head on Visual Representations
The related continual learning methods focus on updating the representation with the backbone model (e.g., [84, 99, 93], and CoDyRA) or newly added representation alignment or prediction head [52, 90]. These two types of methods can be seen as orthogonal. We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in [90].
We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 10. Note that the Dual version of RAIL [90] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 5 and 1 for a fair comparison.